2. Perform a dynamic analysis of the physical system shown in the following illustration. Consider null initial conditions.

a) Define the equations that describe the behavior of each part of the system, in the time domain.
b) Define the equations that describe the behavior of each part of the system, in the frequency domain.
c) Determine the transfer function Θ4 (s) / Τ (s), considering the indicated dynamic parameters.
7) 0,11) (1) 26 N-m-s/rad Ni = 26 N4 = 120 (0) Alo N2 = 110 HN3 = 23 2 N-m/rad
Answer #1

Oz - N; ; 03N4 H E 04 lammating O Ne? 2 Reflect all to middle shaft Te = Ne T = 10-T= 4*2317 Tento rompe De=ing y DP :26 3413